Calculate the pH of a solution that is 0.134 M in benzoic acid (C6H5CO2H) and 0.120 M in hypobromous acid (HBrO).
1
Expert's answer
2020-10-26T14:53:25-0400

The acidic and basic nature of the solution is determined by the pH of a solution. The pH range in the pH scale is 0 to 14. The pH below seven indicates that the solution is acidic. The pH above seven indicates that the solution is basic.


the pH of a solution that is 0.134 M in benzoic acid (C6H5CO2H) 


0.120 M in hypobromous acid (HBrO).


Ka = [H+][B-]/[HB]Benzoic acid dissociates to C6H6COOand H-


this is a 1:1 ratio so the concentrations are equal hence [H+][B-] = [H+]26.3x10-5 = [H+]2/0.16.3x10-6 = [H+]


therefore Hconcentration is 0.0025MpH = -log([H+])


therefore pH = -log(0.0025M) = 2.61